The area of a large rectangular field is 12,750 square yards. The width is 150 feet.
Find the length of the field in yards.
What is the length of the field in feet?
What is the area of the field in square feet?
I know that the equation is A=LW, but I don't know how to convert the feet to yards and then to square yards.
HELP!

Found 2 solutions by scott8148, oberobic:
Answer by scott8148(6628)

1yd=3ft ... 1yd^2=9ft^2 ... width=150ft=50yd ... length=12750/50=255yd=765ft

area=(12750)(9)=114750ft^2

Answer by oberobic(2304)

There are 3 ft in 1 yd, so to convert yards to feet multiply 3. That is 2 yd = 6 ft.
If the width is 150 ft, then divide by 3 to get yds. 150 ft / 3 ft/yd = 50 yd
What is the length? Well, we know the area is 12,750 sq yd. And we know A=W*L, as you said. So 12,750 = 50 * L. Therefore, L = 12,750 / 50 = 2,550 yd (which is a long field indeed.)
To get the length in ft, multiply by 3: 3 ft/yd * 2550 yd = 7650 ft
Finally, 1 sq yd measures 3 ft by 3 ft, so it's 9 sq ft. Given there are 12,750 sq yd, then multiply by 9 to get the sq ft.